OpenStudy (maheshmeghwal9):

Solve for {x}: [x-2/x-4]=[x+4/x-2].

OpenStudy (kaederfds):

x = 5

OpenStudy (maheshmeghwal9):

How?

OpenStudy (kaederfds):

(x-2)(x-2) = (x-4)(x+4)

OpenStudy (maheshmeghwal9):

Ok Thanx!

OpenStudy (anonymous):

i got -3

OpenStudy (maheshmeghwal9):

how?

OpenStudy (anonymous):

hold on lemme double check

OpenStudy (maheshmeghwal9):

k!

OpenStudy (kaederfds):

No no no. The answer is 5 3/1=9/3 3=3 It is true!

OpenStudy (maheshmeghwal9):

ya

OpenStudy (anonymous):

ohh yeah... he is right

OpenStudy (maheshmeghwal9):

i am with kaderfields

OpenStudy (anonymous):

i put wrong number in

OpenStudy (kaederfds):

i am a girl ahaha xD

OpenStudy (anonymous):

haha nice

OpenStudy (maheshmeghwal9):

wow!:)